OpenStudy (anonymous):

Need Calculus help instruction: find the indefinite integral. I'll post the question in a minute.

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ x^2+2x+3 }{ x^3+3x^2+9x } dx\] (There's no interval. you can ignore the question mark.)

OpenStudy (anonymous):

I learned that if the degree of the top one is same or greater than the bottom one, you can do the division, but in this case, the bottom one's larger so I got confused.

OpenStudy (science0229):

or this one, you use the substitution method.

OpenStudy (science0229):

Substitute\[u=x^3+3x^2+9x\]

OpenStudy (science0229):

\[\frac{ 1 }{ 3 }du=(x^2+2x+3)dx\]

OpenStudy (anonymous):

ohh okay, Hold on.. I have to make this clear..

OpenStudy (anonymous):

oh okay I got it Thank you :)

OpenStudy (science0229):

Yep :)

OpenStudy (anonymous):

@science0229 can I ask you another one? of course through my new question page.

OpenStudy (science0229):

Sure.